INAF LBC+LUCIFER queue observing, 2010 December 02 - 03

Observer: K.Boutsia, M. Fumana,V. Testa
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LBC - LUCIFER

Summary

Thin clouds and airplane trails in the sky during twilight - no Sky Flats taken
Seeing <0.8''
RPC error with the trackers, we had to turn OFF - turn ON LBC
Smooth night, overall some thin clouds, great seeing

Data Summary

Project Objects Filter ExpTime Notes
Clementini AndXXI - AndXIX B-bessel, V-bessel 105min + 105min seeing <1.0' - some elongation problems at the beginning
Pentericci cluste1HK 200HK OrderSep 1hr from airmass 1.5 to 3
Cresci HDF 210zJHK+H 1hr25min ARC+FLAT - Tell
Cresci HDF 210zJHK+K 1hr ARC+FLAT - Tell

Details

Begin night with LBC

12:10 Connect LBC to LBT
  • 25Bias_Bino

12:41 - Slew to pointing star WT10_351
  • Pointing correction OK + mirror adjust OK

12:53 - slew to Clementini - AndXXI -dofpia
  • manually give corrections but it converges rather fast

01:24 - observing target Clementini - M31_AndXXI_420s_11
  • Seeing good - <0.8'' from the dimm
  • sky might be still a bit high during the first offset
  • seeing <0.8 in the Blue and <0.6 in the Red
  • blue images slightly elongated...

01:50 - slew to AndXIX - dofpia
  • Blue Rotator has stopped at 41 during slewing for dofpia ... eventually it goes to correct position but it takes some extra time

01:56 - observing target Clementini - M31_AndXIX_420s_11

02:19 - slew to AndXXI - dofpia
  • stopped after 4th iteration - some residual defocusing on the blue side

02:28 - observing target Clementini - M31_AndXXI_420s_12

02:52 - slew to AndXIX - dofpia
  • dofpia,/X2 because we are at very high elevation

03:01 - observing target Clementini - M31_AndXIX_420s_12
  • blue looks better this time

03:26 - Slew to AndXXI - dofpia
  • dofpia,/X2 - it seems to give better results

03:36 - observing target Clementini - M31_AndXXI_420s_13

04:01 - Slew to AndXIX - dofpia
  • dofpia,/X2

04:11 - observing target Clementini - M31_AndXIX_420s_13
  • Blue guide star seemed out focus during first exposure but image looks fine (0.7'')
  • Red image has a slightly worse seeing than blue (0.95'')
  • second exposure both guide stars look fine
  • during last exposure guide star on the red side looked out of focus. Red side seeing worse than on Blue side

04:36 -Slew to AndXXI - dofpia
  • dofpia,/X2

04:47 - observing target Clementini - M31_AndXXI_420s_14

05:14 -Slew to AndXIX - dofpia
  • dofpia

05:17 - observing target Clementini - M31_AndXIX_420s_14
  • a bit of residual astig on the SX

05:40 - problems with the RED tracker during the last exposure
  • RPC host client communication fail
  • Turn LBC OFF- ON
  • last red image not saved!!!
  • SPC not found - reboot please

05:52 - Slew to AndXXI - dofpia
  • stopped after 3rd interaction - some defocus left on the blue side

05:58 - observing target Clementini - M31_AndXXI_420_15

06:23 - slew to AndXIX - dofpia

06:27 - observing target Clementini - M31_AndXIX_420s_15

06:52 - SWITCH TO LUCIFER

07:26 - Pointing correction

07:29 - Slew to Pentericci cluster 1
  • acquisition 0001-0006

07:58 - Pentericci cluster 1 Science 0007-0018
  • 300s mer/normal
  • seeing ~ 0.5" @ z=1.90

09:04 - Calibration Unit IN
  • ARC Ar/Ne 0019-0022
  • increase exp. time to 60s 0023-0026
  • FLAT 0027-0030
  • Calibration unit OUT

09:17 - Slew to Cresci HDFN
  • acquisition 0031-0034

09:37 - Cresci HDFN K_2 0035-0046

10:43 - Calibration unit IN
  • ARC K ==0047-0050==spec_cluster1HK_tslit.txt
  • FLAT K 0051-0056

10:53 - Tell acquisition HDF 0057-0058

10:57 - Tell integration in K for HDF 0059-0066

11:00 - Slew to Cresci HDFN * acquisition 0067-0070

11:13 - Cresci HDFN H_3 0071-0082

12:23 - Lost guiding star - MOS acqusition again 0083-0087

12:30 - Cresci HDFN H_4 0088-0092
  • script was interrupted after the 5th offset because sky was too brightt

12:58 - Calibration Unit IN
  • ARC H 0093-0096
  • FLAT H 0097-0102

13:11 - Slew to Tell acquisition for HDF 0103-0104

13:13 - Sci Int for Tell in H (HDF) 0105-0112

13:16 - CLOSE DOME
  • darks script

-- VincenzoTesta - 03 Dec 2010
Topic revision: r5 - 15 Mar 2012, VincenzoTesta
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback